Find the p and q



1.) x=18;n=100

1
Expert's answer
2022-04-05T12:59:26-0400

p="\frac{x}{n}"

p="\frac{18}{100}"

p=0.18

q=1-p

q=1-0.18

q=0.82

"\frac{}{}"